OpenStudy (anonymous):

which polynomial represents the difference below? 7x8+8x6-2x+9-(4x8+3x7+3)

OpenStudy (aum):

Subtract the like terms. For example, 7x^8 and 4x^8 are like terms.

OpenStudy (aum):

7x^8 + 8x^6 - 2x + 9 - (4x^8 + 3x^7 + 3) = 7x^8 + 8x^6 - 2x + 9 - 4x^8 - 3x^7 - 3 = 7x^8 - 4x^8 - 3x^7 + 8x^6 - 2x + 9 - 3 = ?